# VLM-AD: End-to-End Autonomous Driving through Vision-Language Model Supervision

### v1

@@ 3.3 Auxiliary Loss
Table 2: Ablation study on the contributions of $Q_{1-1}$, $Q_{1-2}$ and $Q_{1-3}$. The best is in bold and the second best is underlined.
4 Experiments
4.1 Settings
Baselines. Our proposed method is a general framework compatible with various end-to-end autonomous driving methods. We validate its effectiveness by applying it to two widely recognized open-source methods, UniAD [18] and VAD [26]. Additionally, we compare it with VLP [40], which projects ego vehicle ground-truth labels into text feature space via CLIP [45] for contrastive learning.
Evaluation Protocol. We focus on the planning task and use standard metrics such as L2 displacement error and collision rate to evaluate performance.
Implementation Details. We use official codes of the UniAD [^1] https://github.com/OpenDriveLab/UniAD and VAD [^2] https://github.com/hustvl/VAD, adhering to the hyperparameters specified in their official implementations. For our proposed VLM-AD, we define two auxiliary task heads, each containing an MHCA block with 8 heads and 3 cross-attention layers, and we set 3 text queries for each of $Q_{1}$ and $Q_{2}$. During training, we set the temperature parameters $\tau_{s}=0.1$ and $\tau_{t}=0.04$ to control the sharpness of the features, and we set $\lambda_{1}=1$ and $\lambda_{2}=0.1$ to balance $\mathcal{L}_{align}$ and $\mathcal{L}_{action}$. All models are trained on 8 NVIDIA H100 GPUs using the PyTorch framework [41]. Complete implementation details, annotation quality analysis, and more experiments are provided in the supplementary material.
4.2 Main Results
Tab. 1 presents the results of applying our proposed VLM-AD to two baselines, UniAD and VAD, as well as comparisons with VLP. Comparing methods ID 0 and 1, we achieve nearly identical planning results using authors’ officially trained checkpoints. For methods IDs 6 and 7, and IDs 12 and 13, we observe some discrepancies between our reproduced results and the reported values, which we attribute to a correction in image configuration in the official codebase [^3] https://github.com/hustvl/VAD/issues/9. From the first section of the table, we observe that VLM-AD significantly outperforms UniAD on both average L2 planning error and average collision rate by introducing $Q_{1}$ and $Q_{2}$. It also surpasses a state-of-the-art baseline VLP in both metrics. Regarding VAD, our VLM-AD consistently outperforms both VAD-Base and VAD-Tiny, particularly on the L2 planning error metric, and it achieves superior performance compared to the VLP in VAD-Base. These results demonstrate the effectiveness and advantages of our VLM-AD approach. Additionally, $Q_{1}$ yields better results than $Q_{2}$, verifying the value of supervising the driving model through rich reasoning information.
4.3 Ablation Study
@@
Sub-Question Contribution. We further analyze the contributions of each sub-question, $Q_{1-1}$, $Q_{1-2}$, and $Q_{1-3}$, within $Q_{1}$. Each sub-question provides specific text information related to the ego vehicle’s current status, predicted future actions, and reasoning. Tab. 2 presents an ablation study of these three sub-questions. The results indicate that each sub-question positively impacts the overall performance, demonstrating that our designed questions provide valuable information for the planning task. Notably, the reasoning feature contributes the most to reducing the L2 planning error, underscoring its importance in enhancing driving performance.
Feature Alignment Loss. We also study alternative options of feature alignment, including minimizing contrastive learning loss in CLIP [45], MSE loss, KL divergence loss [30], or maximizing negative cosine similarity to align the three features from $Q{1}$. The results, shown in Tab. 3, indicate that MSE loss performs slightly better than UniAD, by minimizing the Euclidean distance between features, driving outputs toward their mean that causes information loss during training. Both CLIP loss, KL divergence, and cosine similarity outperform UniAD but are inferior to our alignment loss. This underscores the importance of normalizing with different temperatures to balance the smoothness and sharpness of teacher-student features.
Model Design. We investigate alternative design options in our method. First, we use an MLP layer instead of the MHCA block in our structured action classification head. Second, we study different language models, such as T5 [46], MPNet [50] in addition to CLIP for encoding text annotations from $Q_{1}$. From Tab. 4, we observe that achieves slightly worse L2 performance and the same collision rate. Additionally, we observe that both T5 and MPNet outperform UniAD baseline, but slightly worse than CLIP.
@@
Figure 4: Qualitative comparison between UniAD and our method. The yellow arrow highlights areas where VLM-AD outperforms UniAD. The red box indicates the failure planning command from UniAD, and the purple box represents the predicted three action outputs from our VLM-AD auxiliary text head.
@@
4.4 Visualizations
We provide four visualization examples from the nuScenes dataset, as shown in Fig. 4, to demonstrate the effectiveness of our proposed method. In the first, third, and last row cases, UniAD produces planning trajectories that are winding and lack smoothness, whereas our method generates smoother trajectories that accurately follow the road. Additionally, in the second, third, and last row cases, the commands calculated by the baseline incorrectly suggest a turning intention, while the ego vehicles in these scenarios are actually moving straight. Our action text head correctly outputs the control action “go straight” for all three cases, not only verifying the effectiveness of VLM supervision but also offering interpretability of the model’s decisions.
5 Conclusion
In this work, we presented VLM-AD, a novel approach to enhancing end-to-end autonomous driving models by leveraging vision-language models (VLMs) as auxiliary teachers. By integrating VLM-based annotations through targeted questions that include unstructured reasoning text and structured action labels, we enriched the training process with additional reasoning and action supervision. Our method demonstrated significant improvements in planning accuracy and reduced collision rates on the nuScenes dataset, and offerred interpretability in its output trajectories through action predictions. Importantly, VLM-AD does not require VLMs during inference, making it plug-and-play for real-world deployment without additional inference costs.
Appendix A Implementation Details
When integrating our proposed VLM-AD method into UniAD [18], we follow the joint training protocol defined in UniAD. In the first stage, we initialize the model using BEVFormer [34] weights and train the perception and mapping tasks for 6 epochs. In the second stage, we freeze the image backbone and BEV encoder and perform end-to-end training using our proposed VLM-AD approach for 20 epochs. The model is trained using an initial learning rate of $2\times 10^{-4}$, a learning rate multiplier of 0.1, and the AdamW optimizer [37] with a weight decay of 0.01.
When integrating our proposed VLM-AD method into VAD [26], we adopt the same hyperparameters as in their original implementation. The model is trained using the AdamW optimizer [37] and a cosine annealing scheduler [38], with a weight decay of 0.01 and an initial learning rate of $2\times 10^{-4}$.
Appendix B VLM Annotation
@@
Figure 5: An annotation example of our approach, by projecting the ego vehicle’s future trajectory onto the front-view image.
Figure 6: An annotation example using all 6 images, with ego vehicle’s future trajectory projected onto the front-view image.
B.1 Visual Input
While we use a front-view image (as shown in Fig. 5) as the visual input to the VLM, we also experiment with other alternatives described in Sec. 3, including six images covering the full 360-degree surroundings of the ego vehicle (Fig. 6) and a sequence of consecutive front-view images (Fig. 7). Compared to the first alternative that uses the full 360-degree surrounding as input, our approach yields similar annotations from the VLM while significantly reducing the computational cost by processing a much smaller input image. The second alternative, which utilizes consecutive front-view images, often results in incorrect annotations, such as misidentifying the current action status and failing to detect a left-turn action. This is due to the challenge VLMs face in understanding temporal dynamics, especially from ego-centric visual signals. Furthermore, using consecutive images increases the annotation time by approximately 80% compared to our approach.
@@ B.1 Visual Input
B.2 Annotation Statistics
@@
For structured action annotations using $Q_{2}$, we analyze the distribution of three types of actions. The results are shown in Fig. 8, Fig. 9 and Fig. 10. Approximately 62% of frames are labeled as “go straight”, 89.4% as “no turn action”, and 97.3% as “no lane action”. Notably, no frames are labeled with “reverse” or “turn around”, and only a very small number of frames are labeled as “merge into left lane” or “merge into right lane”. These statistics suggest that the nuScenes dataset has limited diversity in driving actions.
Figure 8: Control action distribution.
Figure 9: Turn action distribution.
Figure 10: Lane action distribution.
@@
B.3 Annotation Quality
@@
B.4 Succesful Annotation Examples
We present three examples to demonstrate the quality of VLM annotations, as shown in Fig. 11, Fig. 12, and Fig. 13.
@@ B.4 Succesful Annotation Examples
In Fig. 11, the VLM accurately identifies the red traffic light and suggests a stop action at the intersection. It also predicts a reasonable future action and clearly explains the rationale behind its decisions.
@@
In Fig. 12, a white van is observed ahead of the ego vehicle but in the opposite lane. The VLM correctly assesses that the van will not affect the ego vehicle’s movement and outputs appropriate driving actions.
@@
In Fig. 13, the ego vehicle is stopped at an intersection on a rainy day. Despite low visibility, the VLM successfully identifies the red traffic light and predicts the future movement based on the traffic light’s status.
Participant Average Score (1-5) Accuracy (%)
1 4.58 4.66 4.26 0.88 0.96 0.98
2 4.34 4.26 4.34 0.86 0.92 0.94
3 4.86 4.66 4.54 0.98 0.84 0.96
4 4.12 4.34 4.40 0.80 0.84 0.94
5 4.50 4.62 4.56 0.98 0.96 0.98
Average 4.48 4.51 4.42 0.90 0.90 0.96
Std 0.28 0.19 0.13 0.08 0.06 0.02
B.5 Imperfect Annotation Examples
We also present three annotation failure cases, illustrated in Fig. 14, Fig. 15, and Fig. 16.
@@ B.5 Imperfect Annotation Examples
In Fig. 14, the VLM accurately recognizes that the traffic light is green and predicts a future right-turn action from its reasoning annotation. However, it incorrectly outputs a left-turn action from the action annotation. Since we query $Q_{1}$ and $Q_{2}$ separately, the response to $Q_{1}$ does not influence $Q_{2}$. One potential solution is to introduce additional prompts to establish a progressive questioning process toward more accurate action annotations.
Figure 11: A successful annotation example where the red traffic light is correctly identified, and the annotations are complete and informative.
Figure 12: A successful annotation example where the white van is correctly identified as irrelevant to the ego vehicle’s driving.
Figure 13: A successful annotation example where the rainy condition and red traffic light are correctly identified, resulting in high-quality annotations.
Figure 14: An imperfect annotation example where reasoning annotations are accurate but the turn action is incorrectly annotated as “turn left”.
Figure 15: An imperfect annotation example where an overly cautious annotation is generated, despite the ego vehicle continuing without stopping or slowing down.
Figure 16: An imperfect annotation example where the pedestrian traffic light is mistakenly considered, leading to an incorrect stop annotation.
@@
In Fig. 15, the VLM outputs “stop” or “move slowly” as the ego vehicle’s current status. While these outputs are plausible, they are inconsistent with the ground truth, as the projected future trajectory indicates that the ego vehicle is currently turning right at the intersection. On the other hand, the action annotation successfully predicts the correct future action.
@@
In Fig. 16, the VLM mistakenly identifies the red pedestrian light as a traffic light and provides incorrect responses.
@@
Overall, despite making occasional mistakes, the VLM demonstrates the ability to generate meaningful annotations that uncover the underlying rationale behind driving decisions, bringing benefits to our proposed method as validated by our experiments. By querying two independent annotation questions, our method is robust to VLM mistakes that often appear in only one of the two responses, as seen in Fig. 14 and Fig. 15. We defer to obtaining more accurate VLM responses as future work to further advance the performance of E2E planning models.
Appendix C Additional Qualitative Examples
We present an enhanced version of Fig. 4 through a set of individual figures: Fig. 17, Fig. 18, Fig. 19, and Fig. 20, where each figure includes the full 6 camera images from the dataset, although our VLM annotation pipeline only focuses on the front-view image, as discussed in Sec. 3.1.
@@ Appendix C Additional Qualitative Examples
In addition to Fig. 4, we provide four extra qualitative examples to compare the planning results of our proposed method against UniAD, as shown in Fig. 21, Fig. 22, Fig. 23, and Fig. 24.
@@
In Fig. 21 and Fig. 22, the planning trajectories generated by UniAD are winding, lack smoothness, and fail to stay within the center of the lane. In contrast, our method produces trajectories that are significantly smoother and stay within the lane boundaries.
@@
Similarly, in Fig. 23 and Fig. 24, the commands generated by UniAD are incorrect, as the ego vehicle is moving forward. However, our action head successfully predicts the correct actions in these scenarios.

### v2

@@ 4.5 Ablation Study
We further analyze the contributions of each sub-question, $Q_{1-1}$, $Q_{1-2}$, and $Q_{1-3}$, within $Q_{1}$. Each sub-question provides specific text information related to the ego vehicle’s current status, predicted future actions, and reasoning. Tab. 3 presents an ablation study of these three sub-questions. The results indicate that each sub-question positively impacts the overall performance, demonstrating that our designed questions provide valuable information for the planning task. Notably, the reasoning feature contributes the most to reducing the L2 planning error, underscoring its importance in enhancing driving performance. We defer additional ablation studies to Sec. C.
5 Conclusion
In this work, we presented VLM-AD, a novel approach to enhancing end-to-end autonomous driving models by leveraging vision-language models (VLMs) as auxiliary teachers. By integrating VLM-based annotations through targeted questions that include unstructured reasoning text and structured action labels, we enriched the training process with additional reasoning and action supervision. Our method significantly improves planning accuracy and collision rates in the nuScenes dataset, while also achieving higher route completion and driving scores under closed-loop evaluation. Importantly, it does not require VLMs at inference time, making it plug-and-play for real-world deployment without additional inference overhead.
Limitations
One limitation of our work is its focus solely on the planning task, with questions centered on behavior and action. In future work, our approach could be extended to generate relevant annotations for perception or prediction, potentially benefiting the entire autonomous driving pipeline. Additionally, we plan to apply our method to more challenging datasets and design a broader set of questions to extract diverse and useful knowledge from VLMs.
Appendix A Related Work
Multi-Task Learning. Multi-task learning (MTL) jointly performs several related tasks using a shared representation through separate branches or heads. This approach leverages shared domain knowledge, enhancing feature robustness and generalization, making it well-suited for end-to-end autonomous driving. In AD systems, auxiliary tasks such as semantic segmentation [54, 55, 56, 57, 58], depth estimation [57, 58], HD mapping, and BEV segmentation [59, 60, 61, 62, 63] are commonly adopted to extract meaningful perception representations for subsequent objects. Beyond vision tasks, prior methods [64, 65] have explored predicting auxiliary signals such as traffic light states and control commands, as well as unifying different trajectory modeling tasks [66, 67], to enhance driving performance. Inspired by the success of multi-task learning, we design novel auxiliary tasks to encourage the model to learn richer feature representations through high-quality reasoning annotations from VLMs, ultimately leading to more reliable planning performance.
Appendix B Implementation Details
When integrating our proposed VLM-AD method into UniAD [2] [^2] https://github.com/OpenDriveLab/UniAD, we follow the joint training protocol defined in UniAD. In the first stage, we initialize the model using BEVFormer [68] weights and train the perception and mapping tasks for 6 epochs. In the second stage, we freeze the image backbone and BEV encoder and perform end-to-end training using our proposed VLM-AD approach for 20 epochs. The model is trained using an initial learning rate of $2\times 10^{-4}$, a learning rate multiplier of 0.1, and the AdamW optimizer [69] with a weight decay of 0.01.
When integrating our proposed VLM-AD method into VAD [3] [^3] https://github.com/hustvl/VAD, we adopt the same hyperparameters as in their original implementation. The model is trained using the AdamW optimizer [69] and a cosine annealing scheduler [70], with a weight decay of 0.01 and an initial learning rate of $2\times 10^{-4}$.
When integrating our proposed VLM-AD method into SparseDrive [32] [^4] https://github.com/swc-17/SparseDrive, we adopt the same hyperparameter settings as in their original implementation. Specifically, we use the SparseDrive-B base model with a ResNet-101 backbone and input image size of 512×1408. The model is trained using the AdamW optimizer [69] and a cosine annealing scheduler [70] in a two-stage training setup, with a learning rate of $3\times 10^{-4}$ and a weight decay of $1\times 10^{-3}$.
For closed-loop evaluation, we integrate our method into the VAD-Base framework and follow the same experimental settings described in their paper [3]. Specifically, as in VAD-Base, the navigation information consists of a sparse goal location and a corresponding discrete navigational command, which are encoded using an MLP and passed to the planning head as part of the input features. Additionally, we retain the original traffic light classification branch used in VAD-Base, which comprises a ResNet-50 backbone followed by an MLP-based classification head. The input to this branch is the upper-middle region of the cropped front-view image, and the resulting feature map is flattened and provided to the planning head to help the model interpret traffic light states.
All experimental settings are kept consistent with the corresponding baseline models, as our proposed method is a plug-and-play module that can be seamlessly integrated into various end-to-end driving frameworks without architectural changes or hyperparameter tuning. This design ensures robustness, cross-model compatibility, and fair comparisons under consistent training conditions, while also highlighting the practicality of VLM-AD for real-world deployment.
@@ Appendix B Implementation Details
Appendix C Ablation Study
Contrastive Learning with Text Annotations. In addition to our proposed auxiliary task, we employ a contrastive learning strategy similar to VLP. We follow VLP [12] to project our generated annotations from $Q_{1}$ using CLIP [19] to obtain three distinct text features. Each text feature is then contrasted with the ego feature individually before we compute the contrastive loss.
The results, presented in Tab. 4, show that contrastive learning improves performance compared to the baseline, though it is still worse than our custom auxiliary tasks. This may be because each frame contains only one ego vehicle, resulting in a single positive sample for contrastive learning that can lead to unstable alignment.
Feature Alignment Loss. We also study alternative options of feature alignment, including minimizing contrastive learning loss in CLIP [19], MSE loss, KL divergence loss [73], or maximizing negative cosine similarity to align the three features from Q1Q{1}. The results, shown in Tab. 4, indicate that MSE loss performs slightly better than UniAD, by minimizing the Euclidean distance between features, driving outputs toward their mean, which causes information loss during training. Both CLIP loss, KL divergence, and cosine similarity outperform UniAD but are inferior to our alignment loss. This underscores the importance of normalizing with different temperatures to balance the smoothness and sharpness of teacher-student features.
Model Design. We investigate alternative design options in our method. First, we use an MLP layer instead of the MHCA block in our structured action classification head. Second, we study different language models, such as T5 [71], MPNet [72] in addition to CLIP for encoding text annotations from $Q_{1}$. From Tab. 5, we observe that MLP achieves slightly worse L2 performance and the same collision rate. Additionally, we observe that both T5 and MPNet outperform the UniAD baseline, but are slightly worse than CLIP.
@@ Appendix C Ablation Study
(a) Annotation using a single front-view image.
(b) Annotation using all 6 images.
Figure 3: Comparison of annotation strategies: (a) only the front-view image is used; (b) all 6 camera views are included, but the trajectory is still projected onto the front view.
Appendix D VLM Annotation
@@
D.1 Visual Input
While we use a front-view image (as shown in Fig. 3(a)) as the visual input to the VLM, we also experiment with other alternatives described in Sec. 3, including six images covering the full 360-degree surroundings of the ego vehicle (Fig. 3(b)) and a sequence of consecutive front-view images (Fig. 4). Compared to the first alternative that uses the full 360-degree surrounding as input, our approach yields similar annotations from the VLM while significantly reducing the computational cost by processing a much smaller input image. The second alternative, which utilizes consecutive front-view images, often results in incorrect annotations, such as misidentifying the current action status and failing to detect a left-turn action. This is due to the challenge VLMs face in understanding temporal dynamics, especially from ego-centric visual signals. Furthermore, using consecutive images increases the annotation time by approximately 80% compared to our approach.
@@ D.1 Visual Input
(a) Control action
(b) Turn action
(c) Lane action
Figure 5: Distributions of control, turn, and lane actions.
D.2 Annotation Statistics
@@
For structured action annotations using $Q_{2}$, we analyze the distribution of three types of actions. The results are shown in Fig. 5(a), Fig. 5(b) and Fig. 5(c). Approximately 62% of frames are labeled as “go straight”, 89.4% as “no turn action”, and 97.3% as “no lane action”. Notably, no frames are labeled with “reverse” or “turn around”, and only a very small number of frames are labeled as “merge into left lane” or “merge into right lane”. These statistics suggest that the nuScenes dataset has limited diversity in driving actions.
@@
Participant Average Score (1-5) Accuracy (%)
1 4.58 4.66 4.26 0.88 0.96 0.98
2 4.34 4.26 4.34 0.86 0.92 0.94
3 4.86 4.66 4.54 0.98 0.84 0.96
4 4.12 4.34 4.40 0.80 0.84 0.94
5 4.50 4.62 4.56 0.98 0.96 0.98
Average 4.48 4.51 4.42 0.90 0.90 0.96
Std 0.28 0.19 0.13 0.08 0.06 0.02
D.3 Annotation Quality
@@
D.4 Succesful Annotation Examples
We present three examples to demonstrate the quality of VLM annotations, as shown in Fig. 6(a), Fig. 6(b), and Fig. 6(c).
@@ D.4 Succesful Annotation Examples
In Fig. 6(a), the VLM accurately identifies the red traffic light and suggests a stop action at the intersection. It also predicts a reasonable future action and clearly explains the rationale behind its decisions.
@@
In Fig. 6(b), a white van is observed ahead of the ego vehicle but in the opposite lane. The VLM correctly assesses that the van will not affect the ego vehicle’s movement and outputs appropriate driving actions.
@@
In Fig. 6(c), the ego vehicle is stopped at an intersection on a rainy day. Despite low visibility, the VLM successfully identifies the red traffic light and predicts the future movement based on the traffic light’s status.
D.5 Imperfect Annotation Examples
We also present three annotation failure cases, illustrated in Fig. 7(a), Fig. 7(b), and Fig. 7(c).
@@ D.5 Imperfect Annotation Examples
In Fig. 7(a), the VLM accurately recognizes that the traffic light is green and predicts a future right-turn action from its reasoning annotation. However, it incorrectly outputs a left-turn action from the action annotation. Since we query $Q_{1}$ and $Q_{2}$ separately, the response to $Q_{1}$ does not influence $Q_{2}$. One potential solution is to introduce additional prompts to establish a progressive questioning process toward more accurate action annotations.
@@
In Fig. 7(b), the VLM outputs “stop” or “move slowly” as the ego vehicle’s current status. While these outputs are plausible, they are inconsistent with the ground truth, as the projected future trajectory indicates that the ego vehicle is currently turning right at the intersection. On the other hand, the action annotation successfully predicts the correct future action.
@@
In Fig. 7(c), the VLM mistakenly identifies the red pedestrian light as a traffic light and provides incorrect responses.
@@
Overall, despite making occasional mistakes, the VLM demonstrates the ability to generate meaningful annotations that uncover the underlying rationale behind driving decisions, bringing benefits to our proposed method as validated by our experiments. By querying two independent annotation questions, our method is robust to VLM mistakes that often appear in only one of the two responses, as seen in Fig. 7(a) and Fig. 7(b). We defer to obtaining more accurate VLM responses as future work to further advance the performance of E2E planning models.
Appendix E Qualitative Examples
@@
We present a total of eight qualitative examples: Fig. 8, Fig. 9, Fig. 10, Fig. 11, Fig. 12, Fig. 13, Fig. 14, and Fig. 15 to compare the planning results of our proposed method against UniAD.
@@
In Fig. 8, Fig. 10, Fig. 11, Fig. 12 and Fig. 13, the planning trajectories generated by UniAD are winding, lack smoothness, and fail to stay within the center of the lane. In contrast, our method produces trajectories that are significantly smoother and stay within the lane boundaries.
@@
Similarly, in Fig. 9, Fig. 10, Fig. 11, Fig. 14, and Fig. 15, the commands generated by UniAD are incorrect, as the ego vehicle is moving forward. However, our action head successfully predicts the correct actions in these scenarios.
@@
(a) A successful annotation example where the red traffic light is correctly identified, and the annotations are complete and informative.
(b) A successful annotation example where the white van is correctly identified as irrelevant to the ego vehicle’s driving.
(c) A successful annotation example where the rainy condition and red traffic light are correctly identified, resulting in high-quality annotations.
Figure 6: Qualitative examples of successful VLM-based annotations.
(a) An imperfect annotation example where reasoning annotations are accurate but the turn action is incorrectly annotated as “turn left”.
(b) An imperfect annotation example where an overly cautious annotation is generated, despite the ego vehicle continuing without stopping or slowing down.
(c) An imperfect annotation example where the pedestrian traffic light is mistakenly considered, leading to an incorrect stop annotation.
Figure 7: Qualitative examples of imperfect VLM-based annotations.